OpenStudy (anonymous):

for the pair of points do the following: (1,4) ; ( 0,1) A. write and equation if the line in point slope form B. write equation in standard form C. write equation in slope intercept form

OpenStudy (fanduekisses):

A. Do you know the point slope form?

OpenStudy (anonymous):

no could you explain it

OpenStudy (fanduekisses):

You will need to know the slope form and the point slope form

OpenStudy (fanduekisses):

Ok so first you need to find the slope.

OpenStudy (anonymous):

i known y = mx + b

OpenStudy (fanduekisses):

\[Slope = \frac{ y _{2} - y_{1}}{ x_{2}-x_{1} }\] so plug in from the points you are given.

OpenStudy (anonymous):

so would it be 1-4 over 0-1?

OpenStudy (fanduekisses):

yes

OpenStudy (anonymous):

then i get -3/-1 which = 3

OpenStudy (fanduekisses):

correct, now for A. you need to write an equation in point slope form Point Slope form = \[y-y_{1}=m(x-x_{1})\] Where m = slope

OpenStudy (anonymous):

would y2 be y or y1?

OpenStudy (fanduekisses):

no, y is just y, keep it as y, do not plug in anything for y or x, ONLY for y1 and x1

OpenStudy (anonymous):

oh okay so it would be y-4 = 3 (x-1)

OpenStudy (fanduekisses):

yes now just distribute the 3 and you're done with part A! :D

OpenStudy (anonymous):

yay now on to part b :)

OpenStudy (fanduekisses):

Ok so part B, Standard form is \[Ax + By = C\]

OpenStudy (fanduekisses):

You use the Point slope form

OpenStudy (anonymous):

how?

OpenStudy (fanduekisses):

You can solve for y

OpenStudy (fanduekisses):

Look, can you see what I did here?

OpenStudy (anonymous):

I dont understand

OpenStudy (fanduekisses):

At first it looks like I was solving for y, then I shifted the y to the other side in order to get it in standard form.

OpenStudy (fanduekisses):

Ok so take the point slope from part A

OpenStudy (anonymous):

oh okay i saw the picture

OpenStudy (fanduekisses):

Do you understand it now?

OpenStudy (anonymous):

yes

OpenStudy (fanduekisses):

Ok so Part c, slope intercept form, this one is easy. y=mx + b

OpenStudy (fanduekisses):

You take the point slope form and solve for y which we already did in step c

OpenStudy (fanduekisses):

I mean in step B*

OpenStudy (anonymous):

y = 3x-1?

OpenStudy (fanduekisses):

Yes woohoo :D

OpenStudy (anonymous):

yayy thank you!

OpenStudy (fanduekisses):

You're very welcome :)

OpenStudy (anonymous):

Do you know how to use matrices?

OpenStudy (fanduekisses):

hmm I'm not quite sure, but let me see the question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Solve each system by using matrices x + 2y - z = -1 x - 4y - z = -4 2x - 2y + 3z = 15

OpenStudy (fanduekisses):

Well, here is a helpful site http://www.mathsisfun.com/algebra/systems-linear-equations-matrices.html

OpenStudy (anonymous):

thank you